### v2

# Extremal Bounds on the Properties of Sigma and Albertson Indices for Non-Decreasing Degree Sequences

Abstract
In this paper, we establish bounds on the topological index — the Sigma index — focusing on analyzing the upper bound of its maximum value, which is known as $\sigma_{\max}(\mathcal{S})=\max\{\sigma(\mathscr{G})\mid\mathscr{G}\in\mathcal{S}\}$, where $\mathcal{S}$ is a class of graphs. We establish precise lower and upper bounds for the Sigma index by leveraging degree sequences $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$, $\mathscr{R}=(t_{1},t_{2},\dots,t_{m})$, and $\mathscr{A}=(a_{1},a_{2},\dots,a_{r})$.
The minimum bound for the Albertson index related to the Sigma index bounds incorporates terms such as $\operatorname{irr}(T)$ and $\left\lfloor\frac{n-2}{a_{r}-t_{m}}\right\rfloor$. These results, rooted in extremal graph theory, enhance the understanding of topological indices in molecular and network analysis.
AMS Classification 2010: 05C05, 05C12, 05C20, 05C25, 05C35, 05C76, 68R10.
Keywords: Trees, Degree sequence, Extremal, Maximum, Irregularity.
1 Introduction
Throughout this paper. Let $G=(V,E)$ be a simple, connected graph, where $n=|V(G)|$, $m=|E(G)|$. Let $\mathscr{D}=(d_{1},d_{2},\dots,d_{n-1},d_{n})$ be a degree sequence where $d_{n}\geqslant d_{n-1}\geqslant\dots\geqslant d_{2}\geqslant d_{1}$, let $\mathscr{R}=(t_{1},t_{2},\dots,t_{m})$ be a degree sequence derived from $\mathscr{D}$ where $t_{1}=(d_{2}-d_{1})/2,t_{2}=(d_{3}-d_{2})/2,\dots,t_{m}=(d_{n}-d_{n-1})/2$. Let $\mathscr{A}=(a_{1},a_{2},\dots,a_{r})$ where $a_{1}=(d_{2}+d_{1})/2,a_{2}=(d_{3}+d_{2})/2,\dots,a_{r}=(d_{n}+d_{n-1})/2$. Let $\lambda_{\mathscr{D}}$ be the average of $\mathscr{D}$, $\lambda_{\mathscr{R}}$ be the average of $\mathscr{R}$ and $\lambda_{\mathscr{A}}$ be the average of $\mathscr{A}$. Hakimi [21] had established the element of $\mathscr{D}$ realizable as a connected graph with $d_{4}\neq 1$. For two vertices $v_{1}$ and $v_{2}$ of $V(G)$ that are categorized into separate classes proved by Schmidt and Druffel [27] according to $\mathscr{D}$, the vertices categorized into separate classes according to their characteristic vector.
D. Vukicevic and B. Furtula [29] had provided results about topological index based on the ratios of geometrical and arithmetical means of end-vertex degrees of edges. Caro and Pepper [11] introduced both upper and lower bounds, connected to indices from $\mathscr{D}$. Akgüneş and Nacaroğlu [8] proved a technique for finding $\mathscr{D}$ where [18] employees $\mathscr{D}$ with variant operators.
The neighborhood [12] of $v\in V(G)$ denote $N(v)$, while consider $N[v]=N(v)\cup\{v\}$ denote the closed neighborhood of $v\in V(G)$. One of the earliest and most influential measures of graph irregularity is the Albertson index, introduced in [4, 24], and defined for a simple graph $G=(V,E)$ as:
$\mathrm{irr}(G)=\sum_{uv\in E(G)}|\deg_{G}(u)-\deg_{G}(v)|.$
This index captures local disparities between adjacent vertex degrees and has inspired numerous variants and generalizations [3, 1]. Gao et al. [17] had presented the impact of topological indices between the chemical characteristics of chemical compounds and drugs and their molecular structures. One of the topological indices which was associated with irregularity in the graph theory called Sigma index, denoted by $\sigma$ and introduced by [19, 5, 24, 10] , which is defined as
$\sigma(G)=\sum_{uv\in E(G)}(\deg_{G}(u)-\deg_{G}(v))^{2}.$
Jahanbanı and Ediz [24] had presented Sigma index of product two graphs $G$ and $H$ as $\sigma(G\times H)=n_{G}\sigma(G)+n_{H}\sigma(H)$, where $n_{G}=|V(G)|$ and $n_{H}=|V(H)|$. In [12] consider $\sigma_{c}(G)$ the sigma chromatic number of $G$, but it is different from Sigma index where it satisfy $1<\sigma_{c}(G)<n$, where the sigma chromatic number it is not only the number of colors being used that is important but the colors themselves.
Gutman et al. [19] proved Sigma index of path $P_{n}$ satisfy $\sigma(P_{n})=2$ and for cycle graph $C_{n}$ in [24] proved $\sigma(C_{n})=0$. Also, for a bipartite graph $K_{n,m}$ provide in [24] Sigma index satisfy $\sigma(K_{n,m})=m(m-n)^{2}$. In 2023, Dimitrov and Stevanović [14] presented the modification of Sigma index called the $\sigma_{t}$ irregularity and defined as
$\sigma_{t}(G)=\sum_{\{u,v\}\subseteq V(G)}(\deg_{G}(u)-\deg_{G}(v))^{2}$
This index was optimised by Filipovski et al. [16] and Alex et al. [7]. There are many studies that have examined the bounds on the Albertson index, some of which have been related to the sigma index, such as [2, 1, 3, 4, 20, 9, 13]. However, with our study of Sigma index on trees through Theorem 2.3, we found that this is not the case satisfied. Therefore, we had to present a study in which we clarify the bounds through Sigma index and highlight some of the results related to the Albertson index. The analysis of graph irregularity indices, such as the Albertson index and the Sigma index, has significantly contributed to our knowledge of graph structures in mathematical chemistry and networks.
This paper is organized as follows. In Section 2, we presented several essential concepts to this paper. In section 3 we provided the impact of bounds among Sigma index. Section 4 we presented the upper bounds of Sigma index.
2 Preliminaries
In this section, we introduce several essential concepts to ensure the reader fully comprehends the results presented in this paper. Specifically, we emphasize that the definition of the Sigma index used here is based on Theorem 2.3 rather than the general concept, and the characterization of its maximum value similarly relies on Theorem 2.3. Through this paper, we point out that all the trees included in the study are caterpillar trees, we consider caterpillar trees denoted by $\mathscr{C}(n,m)$, where $n$ is the number of backbone (or path) vertices and $m$ is the number of pendant vertices attached to each.
Proposition 2.1 ([8]).
Let $\Gamma(G)$ be a monogenic semigroups. Then, the degree sequence given by
$\mathscr{D}(\Gamma(G))=\{1,2,3,\dots,\left\lfloor\frac{n}{2}\right\rfloor-1,\left\lfloor\frac{n}{2}\right\rfloor,\left\lfloor\frac{n}{2}\right\rfloor,\dots,\left\lfloor\frac{n}{2}\right\rfloor+1,\left\lfloor\frac{n}{2}\right\rfloor+2,\dots,n-2,n-1\}.$
Thus, the direct result according to Proposition 2.1 the authors [8] had presented the Albertson index of $\Gamma(G)$ as
$\operatorname{irr}(\Gamma(G))=\begin{cases}\dfrac{n^{3}-4n}{12}&\text{ if }n\text{ even},\\ \dfrac{n^{3}-n}{12}&\text{ if }n\text{ odd}.\end{cases}$
Proposition 2.2 ([30]).
For any graph $G$ with $n$ vertices and $m$ edges. If $\bar{\sigma}(G)=\sigma(\bar{G})$, then the Sigma index satisfy $\sigma(G)+\bar{\sigma}(G)=nM_{1}(G)-4m^{2}$.
Lemma 2.1 ([23]).
Let $T$ be a tree, and let $\mathscr{D}=(d_{1},d_{2},d_{3},d_{4})$ be a degree sequence such that $d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$. Then the Albertson index of the tree $T$ is given by:
$\operatorname{irr}(T)=d_{1}^{2}+d_{n}^{2}+\sum_{i=1}^{4}\lvert d_{i}-d_{i+1}\rvert+\sum_{i=2}^{3}(d_{i}+2)(d_{i}-1)-2.$
Furthermore, to comprehends the result among Lemma 2.1 see [6] where presented graphs with the minimum total irregularity. This foundational work laid the groundwork for subsequent research into graph irregularity measures.
Lemma 2.2 ([26, 15, 25]).
Let $T$ be a tree with $n\geq 1$ and $A=(a_{1},a_{2},\ldots,a_{n})$ be a sequence such that $a_{1}\geq a_{2}\geq\cdots\geq a_{n}$. Then, $(a_{1}+\cdots+a_{n})(a_{1}+a_{n})\geq a_{1}^{2}+\cdots+a_{n}^{2}+na_{1}a_{n},$ and
$n\sum_{i=1}^{n}a_{i}-\left(\sum_{i=1}^{n}\sqrt{a_{i}}\right)^{2}\leq n(n-1)\left(\frac{1}{n}\sum_{i=1}^{n}a_{i}-\left(\prod_{i=1}^{n}a_{i}\right)^{\frac{1}{n}}\right).$ (1)
Theorem 2.3 ([22]).
Let $T$ be a tree of order $n$, and let $\mathscr{D}=(d_{1},\dots,d_{n})$ be a degree sequence such that $d_{n}\geqslant\dots\geqslant d_{1}$. Then, the Sigma index of the tree $T$ is given by:
$\sigma(T)=\sum_{i\in\{1,n\}}(d_{i}+1)(d_{i}-1)^{2}+\sum_{i=2}^{n-1}(d_{i}+2)(d_{i}-1)^{2}+\sum_{i=2}^{n-1}(d_{i}-d_{i+1})^{2}+2n-2.$
Theorem 2.4 ([19]).
Let $\mathcal{S}_{r,k}$ be a duplicate star graph where $\deg_{\mathcal{S}_{r,k}}(u)=k$ and $\deg_{\mathcal{S}_{r,k}}(v)=r$. Then,
$\sigma(\mathcal{S}_{r,k})=(k-1)^{3}+(r-1)^{3}+(k-r)^{2}.$
2.1 Statement Problem
Examining bounds on topological indices is essential in disciplines such as mathematical chemistry and graph theory, as certain indices, including the Wiener index, Randić index, and Zagreb indices, function as numerical representations of graph structures, frequently symbolizing molecular configurations in cheminformatics.
Through this paper, we will answer the following questions:
  1. 1.
    What extent do the bounds affect the Sigma index?
  2. 2.
    How to find the bounds on the Sigma index through the Albertson index?
  3. 3.
    How to graphically represent these bounds so that their actual effect becomes clear to us?
3 Impact Bounds Among Sigma Index
In this section, the fundamental difference between the effects of the Sigma and Albertson indices on connected graphs versus trees is illustrated in Figure 1. The figure clearly highlights that both indices yield small values when applied to connected graphs, whereas the situation changes drastically for trees. According to Theorem 2.3, both indices produce significantly large values on trees, which is further demonstrated through the analysis of their respective upper and lower bounds.
Figure 1: The effects of the Sigma and Albertson indices on connected graphs.
Proposition 3.1.
Let $T$ be a tree of order $n$ with degree sequence $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$. Then the minimum value of the Albertson index satisfies
$0<\frac{2\operatorname{irr}_{\min}}{\Delta(\Delta-1)^{2}}<1.$ (2)
According to Proposition 3.1, we noticed that $2\operatorname{irr}_{\max}<\Delta(\Delta-1)^{2},$ we presented Proposition 3.2 for given the upper bound of the maximum value of Albertson index.
Proposition 3.2.
Let $T$ be a tree of order $n$ with degree sequence $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$. Then the maximum value of the Albertson index satisfies
$\begin{cases}\operatorname{irr}_{\max}>\left\lfloor\frac{2m}{n}\right\rfloor+\left\lceil\frac{2n}{m}\right\rceil+2^{\alpha}&\text{if }d_{n}\leqslant 20,\\ \operatorname{irr}_{\max}<\left\lceil\frac{2n}{m}\right\rceil+2^{\beta}&\text{if }d_{n}>3.\end{cases}$ (3)
Proposition 3.3.
Let $T$ be tree of order $n$, $\Delta_{\mathscr{A}}$ be the maximum degree of $\mathscr{A}$ and $\Delta_{\mathscr{R}}$ be the maximum degree of $\mathscr{R}$. Then, the lower bound of Sigma index satisfy
$\sigma(T)\geqslant\operatorname{irr}(T)+\left\lfloor\frac{n-2}{a_{r}-t_{m}}\right\rfloor+\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}.$ (4)
Proof.
Let $\operatorname{irr}(T)$ be the Albertson index of $T$, $\Delta_{\mathscr{A}}$ be the maximum degree of $\mathscr{A}$ and $\Delta_{\mathscr{R}}$ be the maximum degree of $\mathscr{R}$. Then, clearly $\Delta_{\mathscr{A}}>\Delta_{\mathscr{R}}$ and $\Delta>\Delta_{\mathscr{A}}$. Thus, $(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}<\operatorname{irr}(T)<\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}$. Then, let each leaf $u_{i}$ be adjacent to a vertex $w_{i}$ with degree $\deg_{T}(w_{i})=d_{w_{i}}$. The total contribution from these edges is:
$\sum_{u_{i}\in N(w_{i})}(d_{w_{i}}-1)(d_{w_{i}}-2).$ (5)
Thus, by considering (5) we have
$\sigma(T)-\operatorname{irr}(T)$ $=\sum_{uv\in E(T)}(\deg_{T}(u)-\deg_{T}(v))^{2}-\sum_{uv\in E(T)}|\deg_{T}(u)-\deg_{T}(v)|$
$=\sum_{uv\in E(T)}\left[(\deg_{T}(u)-\deg_{T}(v))^{2}-|\deg_{T}(u)-\deg_{T}(v)|\right]$
$=\sum_{uv\in E(T)}\left[\deg_{T}(u)^{2}+\deg_{T}(v)^{2}-2\deg_{T}(u)\deg_{T}(v)-|\deg_{T}(u)-\deg_{T}(v)|\right]$
$\geqslant\sum_{u\in N(w_{i})}(d_{w_{i}}-1)(d_{w_{i}}-2)+\sum_{v\in N(w_{i})}(d_{w_{i}}-1)(d_{w_{i}}-2).$
Hence, we have
$\sigma(T)-\operatorname{irr}(T)\geqslant\sum_{uv\in E(T)}\left((\deg_{T}(u)-1)(\deg_{T}(u)-2)+(\deg_{T}(v)-1)(\deg_{T}(v)-2)\right).$ (6)
On the other hand, for the second term, since we observe that $n-2>a_{r}-t_{m}$, and to achieve this difference of $a_{r}-t_{m}<\Delta$, we obtain
$\sigma(T)\geqslant\operatorname{irr}(T)+\left\lfloor\frac{n-2}{a_{r}-t_{m}}\right\rfloor.$ (7)
From relations (6) and (7), considering the quantity $\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}$ as a constant term, where $(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}<\operatorname{irr}(T)<\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}$.Thus, the relation (4) remains valid by adding this quantity to relation (7). ∎
The upper bound of Sigma index holds among Proposition 3.4 by considering the maximum degree of $\mathscr{A}$ and the maximum degree of $\mathscr{R}$.
Proposition 3.4.
Let $T$ be tree of order $n$, $\Delta_{\mathscr{A}}$ be the maximum degree of $\mathscr{A}$ and $\Delta_{\mathscr{R}}$ be the maximum degree of $\mathscr{R}$. Then, the upper bound of Sigma index satisfy
$\sigma(T)\leqslant\sum_{v\in V(T)}\deg_{T}(v)^{3}+\operatorname{irr}(T)+\left\lfloor\frac{n-2}{a_{r}-t_{m}}\right\rfloor+\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}.$ (8)
Proof.
According to Proposition 3.3 we find that $(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}<\operatorname{irr}(T)<\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}$. Thus,
$\sigma(T)\leqslant\sum_{v\in V(T)}\deg_{T}(v)^{3}+\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}.$ (9)
Then, by considering $n<\Delta(\Delta_{\mathscr{A}}-\Delta_{\mathscr{R}})^{2}$ we have $\sigma(T)\leqslant(n-\Delta_{\mathscr{A}})^{2}+(n-\Delta_{\mathscr{R}})^{2}$. Thus, the Albertson index satisfy $\operatorname{irr}(T)\geqslant 2(\Delta_{\mathscr{A}}^{2}-\Delta_{\mathscr{A}}\Delta_{\mathscr{R}}+\Delta_{\mathscr{R}}^{2})$ . Then,
$\sigma(T)\leqslant 2n^{2}-2(\Delta_{\mathscr{A}}\Delta_{\mathscr{R}})n+2(\Delta_{\mathscr{A}}^{2}-\Delta_{\mathscr{A}}\Delta_{\mathscr{R}}+\Delta_{\mathscr{R}}^{2}).$ (10)
Therefore, the relationship between Sigma index and the Albertson index holds $\sigma(T)\geqslant\Delta_{\mathscr{A}}\operatorname{irr}(T)+2(\Delta_{\mathscr{A}}\Delta_{\mathscr{R}})n$. Then,
$\sigma(T)\leqslant 2n^{2}+\sum_{v\in V(T)}\deg_{T}(v)^{3}+\operatorname{irr}(T),$ (11)
from (9) and (11) yields
$\sigma(T)\leqslant\sum_{v\in V(T)}\deg_{T}(v)^{3}+\operatorname{irr}(T)+\left\lfloor\frac{n-2}{a_{r}-t_{m}}\right\rfloor.$ (12)
Finally, from (11) and (12) we find that the upper bound of Sigma index satisfy by (8). ∎
3.1 Study Bounds of Sigma Index Among Albertson Index
The following Lemma 3.1 states that, given specific constraints on the degree sequence $\mathscr{D}$ and order of a graph, the Sigma index exceeds the irregularity index by a strictly positive quantity consisting of explicit additive components with condition $n\leqslant\Delta_{\mathscr{A}}(a_{r}-a_{1})+\Delta_{\mathscr{R}}(t_{m}-t_{1})<\operatorname{irr}(T)$.
Lemma 3.1.
Let $T$ be tree of order $n$, $\Delta_{\mathscr{A}}$ be the maximum degree of $\mathscr{A}$ and $\Delta_{\mathscr{R}}$ be the maximum degree of $\mathscr{R}$. If $n\leqslant\Delta_{\mathscr{A}}(a_{r}-a_{1})+\Delta_{\mathscr{R}}(t_{m}-t_{1})<\operatorname{irr}(T)$. Then, the lower bound of Sigma index satisfy
$\sigma(T)\geqslant\operatorname{irr}(T)+\dfrac{\left\lfloor\dfrac{2n}{a_{r}-a_{1}}\right\rfloor+\left\lceil\dfrac{2m}{n}\right\rceil}{n}+4n\Delta.$ (13)
Proof.
Recall $\mathscr{D}=(d_{1},d_{2},\dots,d_{n-1},d_{n})$ be a degree sequence where $d_{n}\geqslant d_{n-1}\geqslant\dots\geqslant d_{2}\geqslant d_{1}$, and $\mathscr{A}=(a_{1},a_{2},\dots,a_{r})$ where $a_{1}=(d_{2}+d_{1})/2,a_{2}=(d_{3}+d_{2})/2,\dots,a_{r}=(d_{n}+d_{n-1})/2$. Then, we noticed that $\Delta>a_{r}-a_{1}$ and $\sigma(T)>\operatorname{irr}(T)+\Delta_{\mathscr{A}}(a_{r}-a_{1})$. Then, we consider
$\operatorname{irr}(T)\leqslant\left\lfloor\frac{2n^{2}}{a_{r}-a_{1}}\right\rfloor.$ (14)
The study of this bound through the approach related to the Albertson index establishes the relationship we aim to reach, noting that through relation (14) we have obtained the values for the bound $2m/n<\Delta$. Therefore, according to Theorem 2.3 we find that
$\sigma(T)-\operatorname{irr}(T)\geqslant\left\lfloor\frac{2n^{2}}{a_{r}-a_{1}}\right\rfloor+4\left\lceil\dfrac{2m}{n}\right\rceil\Delta.$ (15)
Thus, by considering $2m/n<n$ and $4n>\Delta$ we find that $\sigma(T)<4n^{2}\Delta$. Thus, according to (15) noticed that $\Delta>a_{r}-a_{1}$ and $n>a_{r}-a_{1}$. Then,
$\sigma(T)\geqslant\operatorname{irr}(T)+\left\lfloor\frac{2n^{2}}{a_{r}-a_{1}}\right\rfloor+\frac{4n^{2}\Delta}{a_{r}-a_{1}}.$ (16)
These upper bounds for the Sigma index indeed satisfy relation (16). Moreover, using relation (15) to confirm that
$\sigma(T)\geqslant\dfrac{\left\lfloor\dfrac{2n}{a_{r}-a_{1}}\right\rfloor+\left\lceil\dfrac{2m}{n}\right\rceil}{n}+4n\Delta,\quad\operatorname{irr}(T)\geqslant\frac{n}{a_{r}-a_{1}}+n,$
along with relation (14) and (16), relation (13) is fulfilled, which results in the lower bound of the Sigma index. ∎
Lemma 3.1 frequently arises in the context of extremal graph theory and the analysis of topological indices (the Albertson and Sigma indices). This lemma is often employed to establish inequalities (as we show that among Figure 1) or bounds in Proposition 3.3 and 3.4 on these indices by utilizing degree sequences and the maximum vertex degree.
Lemma 3.2.
Let $T$ be tree of order $n$, $\Delta_{\mathscr{A}}$ be the maximum degree of $\mathscr{A}$ and $\Delta_{\mathscr{R}}$ be the maximum degree of $\mathscr{R}$. Then, the lower bound of Sigma index satisfy
$\sigma(T)\geqslant\sqrt{\lambda_{\mathscr{D}}\sum_{i=1}^{n}d_{i}^{3}}-\left(\left\lfloor\frac{2n}{\lambda_{\mathscr{A}}}\right\rfloor+\left\lceil\frac{2m}{\lambda_{\mathscr{R}}}\right\rceil\right)+(n-\Delta)^{2}.$ (17)
Proof.
Recall $\lambda_{\mathscr{D}}$ be the average of $\mathscr{D}$, $\lambda_{\mathscr{R}}$ be the average of $\mathscr{R}$ and $\lambda_{\mathscr{A}}$ be the average of $\mathscr{A}$. According to Proposition 3.4 we find that
$\sigma(T)\leqslant 2n^{2}+\sum_{v\in V(T)}\deg_{T}(v)^{3}.$
Thus, when $\lambda_{\mathscr{D}}<n$ and $\lambda_{\mathscr{D}}\leqslant\Delta$. Also, for $\lambda_{\mathscr{A}}$ and $\lambda_{\mathscr{R}}$ satisfied. Thus, we need to compare the tem $2n/\lambda_{\mathscr{A}}$ and the term $2m/\lambda_{\mathscr{R}}$ with the term $2n/\lambda_{\mathscr{D}}$ by considering the term $2n/\lambda_{\mathscr{D}}$ is constant term (by examining values we find that the term equals 14). Thus,
$\frac{2n}{\lambda_{\mathscr{D}}}<\left\lfloor\frac{2n}{\lambda_{\mathscr{A}}}\right\rfloor+\left\lceil\frac{2m}{\lambda_{\mathscr{R}}}\right\rceil.$ (18)
Therefore, from (18) by considering $(n-\Delta)^{2}$ is growing up and for Sigma index satisfy $\sigma(T)>(n-\Delta)^{2}$. Then, we find that $\sigma(T)>(n-\Delta)^{2}+\frac{2n}{\lambda_{\mathscr{D}}}$. Thus, we have
$\sigma(T)\geqslant(n-\Delta)^{2}+\frac{2n}{\lambda_{\mathscr{D}}}+\left\lfloor\frac{2n}{\lambda_{\mathscr{A}}}\right\rfloor+\left\lceil\frac{2m}{\lambda_{\mathscr{R}}}\right\rceil.$ (19)
Then, clearly $2n^{2}>\lambda_{\mathscr{D}}$ and $\sqrt{\lambda_{\mathscr{D}}n^{3}}\leqslant(n-\Delta)^{2}$. Thus, for $1<b\leqslant 2n/\lambda_{\mathscr{D}}$ we find that $\sigma(T)\geqslant\sqrt{\lambda_{\mathscr{D}}bn^{3}}$. Thus, $\sigma(T)\leqslant(n-\Delta)^{2}+\sqrt{2n^{4}}.$ Hence, we have
$\sigma(T)\geqslant(n-\Delta)^{2}+\sqrt{\lambda_{\mathscr{D}}\sum_{i=1}^{n}d_{i}^{3}}.$ (20)
This relation holds when considering the difference between terms with the constant term satisfy
$\left(\left\lfloor\frac{2n}{\lambda_{\mathscr{A}}}\right\rfloor+\left\lceil\frac{2m}{\lambda_{\mathscr{R}}}\right\rceil\right)-\frac{2n}{\lambda_{\mathscr{D}}}\leqslant\sqrt{\lambda_{\mathscr{D}}\sigma(T)}.$
Finally, we find that from (18)–(20) the relationship (17) holds. ∎
Among Lemma 3.2, we provide the lower bound combines the terms of both $\lambda_{\mathscr{D}}$, $\lambda_{\mathscr{A}}$ and $\lambda_{\mathscr{R}}$, where $2n^{2}>\lambda_{\mathscr{D}}$ and $\sqrt{\lambda_{\mathscr{D}}n^{3}}\leqslant(n-\Delta)^{2}$. Lemma 3.2 emphasizes a stronger dependence on $\lambda_{\mathscr{D}}$ where $\left\lfloor 3\Delta t/2\right\rfloor\leqslant 3t(n-1)/2$ by considering the Albertson index $\operatorname{irr}(T)$ to measure degree heterogeneity.
Lemma 3.3.
Let $T$ be tree of order $n$, $\Delta_{\mathscr{A}}$ be the maximum degree of $\mathscr{A}$ and $\Delta_{\mathscr{R}}$ be the maximum degree of $\mathscr{R}$. Then, the lower bound of Sigma index satisfy
$\sigma(T)\geqslant\frac{1}{3}\lambda_{\mathscr{D}}^{2}\left(\left\lfloor\dfrac{3n+1}{2}\right\rfloor+\left\lceil\dfrac{3m+1}{2}\right\rceil+\left\lfloor\dfrac{3\Delta+2n}{4}\right\rfloor\right)-\sum_{v\in V(T)}\deg_{T}(v)^{3}+\operatorname{irr}(T).$ (21)
Proof.
Assume $\mathscr{D}=(d_{1},d_{2},\dots,d_{n-1},d_{n})$ be a degree sequence where $d_{n}\geqslant d_{n-1}\geqslant\dots\geqslant d_{2}\geqslant d_{1}$. We have previously found that the Sigma index attains the lower bound with respect to the sum of the cubes of the degrees, and the same applies to the Albertson index. Therefore, we must prove the first part of inequality (21). Thus, let us assume $t>2,t\in\mathbb{N}$, and then we need to prove the validity of the following statement
$t\left\lfloor\dfrac{3n+1}{2}\right\rfloor+2t\left\lceil\dfrac{3m+1}{2}\right\rceil+\left\lfloor\dfrac{3\Delta t}{2}\right\rfloor\leqslant\sigma(T).$ (22)
The maximum degree $\Delta$ satisfies $1\leq\Delta\leq n-1$, where $\Delta=n-1$ occurs in a star graph, and $\Delta=2$ in a path graph with $n\geq 3$. By considering m=n1m=n-1, we find that the relationship (22) satisfy
$t\left\lfloor\dfrac{3n+1}{2}\right\rfloor+2t\left\lceil\dfrac{3n-2}{2}\right\rceil+\left\lfloor\dfrac{3\Delta t}{2}\right\rfloor\leqslant\sigma(T).$ (23)
Case 1:
If $n$ is even, where $n=2k$. Then, from (23) should be prove the lower bound of Sigma index as
$t\left\lfloor\dfrac{6k+1}{2}\right\rfloor+2t\left\lceil 3k-1\right\rceil+\left\lfloor\dfrac{3\Delta t}{2}\right\rfloor\leqslant\sigma(T).$ (24)
Then, the term $2t$ is even, the inequality (24) holds
$t\left\lfloor\dfrac{6k+1}{2}\right\rfloor+2t\left\lceil 3k-1\right\rceil+\left\lfloor\dfrac{3\Delta t}{2}\right\rfloor=9tk-2t+\left\lfloor\dfrac{3\Delta t}{2}\right\rfloor.$ (25)
Thus, for the term of the maximum degree $\Delta$ satisfy $\left\lfloor 3\Delta t/2\right\rfloor\leqslant 3t(n-1)/2$. Then, from (24) and (25) we have
$9tk-2t+\left\lfloor\dfrac{3\Delta t}{2}\right\rfloor\leqslant 12tk-2t+\frac{3t}{2}.$ (26)
Therefore, according to (26) for the term $t/2\leqslant k$. Then, $\sigma(T)\geqslant 12tk-t/2$. Thus, the relationship (24) holds.
Case 2:
If $n$ is odd, where $n=2k+1$. Then, from (23) should be prove the lower bound of Sigma index as
$t\left\lfloor 3k+2\right\rfloor+2t\left\lceil\dfrac{6k+1}{2}\right\rceil+\left\lfloor\dfrac{3\Delta t}{2}\right\rfloor\leqslant\sigma(T).$ (27)
Then,
$t\left\lfloor 3k+2\right\rfloor+2t\left\lceil\dfrac{6k+1}{2}\right\rceil+\left\lfloor\dfrac{3\Delta t}{2}\right\rfloor=9tk+4t+\left\lfloor\dfrac{3\Delta t}{2}\right\rfloor$ (28)
Thus, when $\left\lfloor 3\Delta t/2\right\rfloor\leqslant 3t(n-1)/2$, from (28) satisfy
$9tk+4t+\left\lfloor\dfrac{3\Delta t}{2}\right\rfloor\leqslant 12tk+4t.$ (29)
Thus, according to (29) for the term $4t\leqslant k$ we find that (27) holds. Now, from Case 1 and Case 2 we find that (23) holds. Thus, we noticed that
$\sigma(T)\leqslant\frac{1}{3}\lambda_{\mathscr{D}}^{2}\left(\left\lfloor\dfrac{3n+1}{2}\right\rfloor+\left\lceil\dfrac{3m+1}{2}\right\rceil+\left\lfloor\dfrac{3\Delta+2n}{4}\right\rfloor\right).$ (30)
Therefore, Sigma index satisfy $\sigma(T)\leqslant\operatorname{irr}(T)+n^{3}$ and when $n^{3}\leqslant\sum_{v\in V(T)}\deg_{T}(v)^{3}$ we find that (21) holds and satisfy the lower bound of Sigma index. ∎ Furthermore, we presented the following Table 1, to emphasize that results among Lemma 3.3, where $T_{1}$ and $T_{2}$ given as
$T_{1}=\left\lfloor\dfrac{3n+1}{2}\right\rfloor+\left\lceil\dfrac{3m+1}{2}\right\rceil+\left\lfloor\dfrac{3\Delta+2n}{4}\right\rfloor,\quad T_{2}=\lambda_{\mathscr{D}}^{2}T_{1}.$
Degree Sequence $T_{1}$ $T_{2}$ $\operatorname{irr}(T)$ $\sigma(T)$
(3,5,7,5,6,8,10) 160 2107 260 2248
(7,8,10,11,12,14,15) 280 11293 810 10747
(11,11,13,17,18,20,20) 399 32842 1694 31070
(15,14,16,23,24,26,25) 519 72197 2912 68563
(19,17,19,29,30,32,30) 637 134229 4464 128572
(23,20,22,35,36,38,35) 757 224942 6350 216443
(27,23,25,41,42,44,40) 876 348993 8570 337522
(31,26,28,47,48,50,45) 996 512397 11124 497155
Table 1: Sigma index attains the lower bound among Lemma 3.3.
The positions labeled $n,\sigma,\operatorname{irr}(T),T_{1}$ and $T_{2}$ correspond to the variables whose pairwise correlation coefficients are displayed in the correlation matrix as
$M_{\mathscr{D}}=\begin{pmatrix}1.000000&0.929672&0.974594&0.999999&0.894803\\ 0.929672&1.000000&0.987695&0.929695&0.993202\\ 0.974594&0.987695&1.000000&0.974594&0.966635\\ 0.999999&0.929695&0.974594&1.000000&0.894845\\ 0.894803&0.993202&0.966635&0.894845&1.000000\end{pmatrix}$
Noticed that, regression coefficients [94764.10761811, 27387.84112146][\,-94764.10761811\,,\,27387.84112146\,], intercept 506577.67445476603-506577.67445476603, Model $R^{2}$ is $0.864810869832136$ and predicted Sigma for $n=350$ is 32304623.284719232-32304623.284719232.
Through the following Figure 2, we present analyze the data among Table 1 to consider the results in Lemma 3.3 cleared.
Figure 2: Analysis of data among Table 1.
4 Bounds of Sigma Index
Let $\mathcal{S}$ be a class of graphs. Then, we have $\sigma_{\max},\sigma_{\min}$ of a graph $G$, where:
$\sigma_{\max}(\mathcal{S})=\max\{\sigma(\mathscr{G})\mid\mathscr{G}\in\mathcal{S}\},$
$\sigma_{\min}(\mathcal{S})=\min\{\sigma(\mathscr{G})\mid\mathscr{G}\in\mathcal{S}\}.$
For the purposes of this study, by Proposition 4.1 we presented the lower bound of Sigma index. Since $T$ is a tree with $n$ vertices, it has exactly n1n-1 edges, and the sum of degrees is 2(n1)2(n-1).
Proposition 4.1.
Let $T$ be a tree of order $n$ with the maximum degree $\Delta$, let $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be a degree sequence where $d_{n}\geqslant d_{n-1}\geqslant\dots\geqslant d_{2}\geqslant d_{1}$. Then, the lower bound of Sigma index satisfy
$\sigma(T)>\frac{1}{2\lambda}\left(n^{3}+n+\Delta(\Delta-1)^{2}\right).$ (31)
Lemma 4.1 provides an important upper bound on the Sigma index of a tree graph.
Lemma 4.1.
Let $T$ be a tree with $n$ vertices, $m$ edges, maximum degree $\Delta$, and the Albertson index $\operatorname{irr}(T)$, let $p$ be the prime number. Then, the following upper bound holds for the Sigma index:
$\sigma(T)\leqslant 2^{p}(\operatorname{irr}(T)+2m)+\Delta(\Delta-1)^{2}.$ (32)
4.1 The Upper Bound of Sigma Index
Furthermore, among the maximum value of Sigma index, we present the upper bound of $\sigma_{\max}$ in Theorem 4.2 by considering $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be non-decreasing degree sequence.
Theorem 4.2.
Let $T$ be tree of order $n$, let $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be non-decreasing degree sequence. Then, the upper bound of maximum value of Sigma index satisfy
$\sigma_{\max}(T)\leqslant\dfrac{1}{2(\Delta-3)}\left(\left\lfloor\dfrac{3n^{2}}{4}\right\rfloor\left\lceil\dfrac{n^{2}}{4}\right\rceil\right).$ (33)
Proof.
Assume $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be non-decreasing degree sequence. Then, let discuss the equation (33) as two parts, the first part related to $\Delta$ and the second part related to $3n^{2}/4$ and $n^{2}/4$. Thus, we need to prove the following relationship
$\sigma_{\max}(T)\leqslant t\left\lfloor\dfrac{3n^{2}}{4}\right\rfloor+\frac{1}{t}\left\lceil\dfrac{n^{2}}{4}\right\rceil,$ (34)
where $t>2$ and $t\in\mathbb{N}$. Since both floor and ceiling functions round to integers. Then, we have to emphasize that $3n^{2}/4$ and $n^{2}/4$,
$n^{2}=\frac{3n^{2}}{4}+\frac{n^{2}}{4}.$ (35)
Hence, more precisely, let $\delta$ be the minimum degree of $T$, from (35) the relationship (34) holds
$\left\lfloor\dfrac{3n^{2}}{4}\right\rfloor+\left\lceil\dfrac{n^{2}}{4}\right\rceil=n^{2}+\delta.$ (36)
Then, we discuss the following cases according to whether the values of the vertices are odd or even to clarify these bounds with respect to the value of $t$.
Case 1:
If $n$ is even, where $n=2k$. Then, to prove that (34) we need to prove the following relationship
$\sigma_{\max}(T)\leqslant t\left\lfloor\dfrac{3(2k)^{2}}{4}\right\rfloor+\frac{1}{t}\left\lceil\dfrac{(2k)^{2}}{4}\right\rceil.$ (37)
Thus,
$t\left\lfloor\dfrac{3(2k)^{2}}{4}\right\rfloor+\frac{1}{t}\left\lceil\dfrac{(2k)^{2}}{4}\right\rceil\leqslant 3tk^{2}+\frac{k^{2}}{2t},$ (38)
Therefore, when $t>2$ we find that $\sigma_{\max}(T)\leqslant(6(tk)^{2}+k^{2})/2t+2\delta t$. Then, the relationship (37) holds. Thus, when $n=2k$ we find that (34) holds.
Case 2:
If $n$ is odd, where $n=2k+1$. Then, to prove that (34) we need to prove the following relationship
$\sigma_{\max}(T)\leqslant t\left\lfloor\dfrac{3(2k+1)^{2}}{4}\right\rfloor+\frac{1}{t}\left\lceil\dfrac{(2k+1)^{2}}{4}\right\rceil.$ (39)
Thus,
$t\left\lfloor\dfrac{3(2k+1)^{2}}{4}\right\rfloor+\frac{1}{t}\left\lceil\dfrac{(2k+1)^{2}}{4}\right\rceil\leqslant t\frac{12k^{2}+12k+3}{4}+\frac{4k^{2}+4k+1}{4t},$ (40)
We find in this case a discussion different from what was communicated in Case 1. Therefore, for these values regarding $t>2$, we notice that
$t\left\lfloor\dfrac{3(2k+1)^{2}}{4}\right\rfloor+\frac{1}{t}\left\lceil\dfrac{(2k+1)^{2}}{4}\right\rceil\leqslant 3t(k^{2}+k+1)+\frac{k^{2}+k}{t}.$
Therefore, from (40) we notice that $\sigma_{\max}(T)\leqslant 3t(k^{2}+k+1)+\frac{k^{2}+k}{t}+2t\delta+1$. Thus, the relationship (39) hods.
Now, by considering $n>2(\Delta-3)$ and according to (37) and (39) we emphasize that $t\leq n^{2}+\delta$. Thus, the upper bound of maximum Sigma index satisfy,
$\sigma_{\max}(T)\leqslant\left(\left\lfloor\dfrac{3n^{2}}{4}\right\rfloor\left\lceil\dfrac{n^{2}}{4}\right\rceil\right).$ (41)
Thus, from (41) for the term $\frac{1}{2(\Delta-3)}$ satisfy $4\leq\Delta-3\leqslant n/4$. The relationship (33) holds. ∎ In particular, the previous theorem offers a sharp bound for the maximum bound of the Sigma index. Furthermore, let $\eta>0$ defined by
$\eta=\left\lceil\frac{2n\Delta}{m}\right\rceil.$
Theorem 4.3.
Let $T$ be tree of order $n$, let $\eta$ be an integer, $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be non-decreasing degree sequence with $\lambda_{\mathscr{D}}$ the average of elements $\mathscr{D}$. Then, the upper bound of Sigma index satisfy
$\sigma(T)\leqslant\left\lfloor\frac{2n^{2}}{3\lambda_{\mathscr{D}}}\right\rfloor+\dfrac{2^{\eta}(m-\Delta)^{2}}{5(n-1)^{3}}.$ (42)
Proof.
Recall $\eta>0$ and a non-decreasing degree sequence $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ with $\lambda_{\mathscr{D}}$ the average of elements $\mathscr{D}$. Then, according to the relationship of $k$ we consider that
$\frac{n}{3}<\eta<\left\lfloor\frac{2n^{2}}{3\lambda_{\mathscr{D}}}\right\rfloor.$ (43)
Thus, according to the value of $\eta$ yields $\sigma(T)>\eta$. Then, it holds $\sigma(T)>\left\lfloor 3n^{2}/3\lambda_{\mathscr{D}}\right\rfloor$. Thus, while $\eta<(\Delta-1)^{2}$ holds $\sigma(T)<\eta(\eta-1)^{2}$. Then, the Sigma index achieves the upper bound according to the following relation, by considering the significant difference between $\Delta$ and $\eta$ as
$\sigma(T)\leqslant\left\lfloor\frac{2n^{2}}{3\lambda_{\mathscr{D}}}\right\rfloor+\eta(\eta-1)^{2}.$ (44)
Through this relation, we show that when $\Delta<\eta<n$ and $\eta<3\lambda_{\mathscr{D}}$, then $\sigma(T)<2^{\eta}$ is satisfied, and accordingly, we obtain $2^{\eta}>\eta(\eta-1)^{2}$. Hence, the following relation is satisfied
$\sigma(T)\geqslant\frac{2^{\eta}}{\eta(\eta-1)^{2}}.$ (45)
Actually, relation (45) holds directly to this ratio has been achieved for $\eta$. Therefore, for the bound related to quantity $2^{\eta}(m-\Delta)^{2}$, it clearly follows that relation $\sigma(T)<2^{\eta}(m-\Delta)^{2}$ holds. Thus, according to (44) and (45) we have
$\sigma(T)\leqslant\frac{2^{\eta}(m-\Delta)^{2}}{n\eta(\eta-1)^{2}}.$ (46)
Therefore, by considering $\eta(\eta-1)^{2}<(n-1)^{2}$ we find that $\sigma(T)<2^{\eta}(m-\Delta)^{2}/(n-1)^{2}$. To obtain a smaller term so that the difference between the upper bound and the Sigma index is sufficiently small. From (45) and (46) we find that (42) holds. ∎
Based on the relations discussed in the proof of Theorem 4.3 and according to [28], particularly relations (44) and (46), we directly present in equation (47) a clear and explicit result that provides a sharp lower bound for the Sigma index.
Corollary 4.4.
Let $T$ be tree of order $n$, let $\eta$ be an integer, $\operatorname{irr}(T)$ be the Albertson index and $\lambda_{\mathscr{D}}$ the average of elements $\mathscr{D}$. Then,
$\sigma(T)>4n-2\eta\lambda_{\mathscr{D}}-(n-\eta)\left\lfloor\frac{n}{n-\eta}\right\rfloor^{2}+(n-\eta)\left\lfloor\frac{n}{n-\lambda_{\mathscr{D}}}\right\rfloor.$ (47)
Let $\eta_{1}$ be a new parameter related to $\eta$ where $2<\eta_{1}\leqslant 4$.
Theorem 4.5.
Let $T$ be tree of order $n$, let $\eta,\eta_{1}$ be an integers, $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be non-decreasing degree sequence with $\lambda_{\mathscr{D}}$ the average of elements $\mathscr{D}$. Then, the upper bound of Sigma index satisfy
$\sigma(T)\leqslant\eta_{1}\left\lfloor\frac{n}{n-\eta}\right\rfloor+\eta_{1}\left\lceil\frac{n}{\eta-\lambda_{\mathscr{D}}}\right\rceil+\sum_{v\in V(T)}\deg_{T}(v)^{3}.$ (48)
Proof.
Recall $\eta_{1}$ be a new parameter related to $\eta$ where $2<\eta_{1}\leqslant 4$ and we will consider $\eta_{1}=2^{n}/(n-\eta)!$ and $\delta$ be the minimum degree of $T$. Assume $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be non-decreasing degree sequence with $\lambda_{\mathscr{D}}$ the average of elements $\mathscr{D}$. Then, by considering Corollary (4.4), we need to prove the following relationship
$\sigma(T)\leqslant t\eta_{1}\left\lfloor\frac{n}{n-\eta}\right\rfloor+\frac{1}{t}\eta_{1}\left\lceil\frac{n}{\eta-\lambda_{\mathscr{D}}}\right\rceil,$ (49)
where $t>2$ and $t\in\mathbb{N}$. Thus, we will discuss the following cases.
Case 1:
If $n$ is even, where $n=2k$. Then, we have according to (49) the relationship should be
$\sigma(T)\geqslant t\eta_{1}\left\lfloor\frac{2k}{2k-\eta}\right\rfloor+\frac{1}{t}\eta_{1}\left\lceil\frac{2k}{\eta-\lambda_{\mathscr{D}}}\right\rceil,$ (50)
Then, we noticed that $t\leqslant 2k/(2k-\eta)$ and $t\leqslant 2k/(\eta-\lambda_{\mathscr{D}})$. Then, we conclude that
$t\leq\min\left(\frac{2k}{2k-\eta},\frac{2k}{\eta-\lambda_{\mathscr{D}}}\right).$
Therefore, according to the value of $\eta_{1}$ we find that $t\eta_{1}\leqslant 2k$ and $2k/(2k-\eta)\geqslant 1$. Then, to determine the approximate values of $t$, we observe from $\eta$ and $\eta_{1}$ that they satisfy the following inequality.
$\eta_{1}<\left\lceil\frac{2k}{\eta-\lambda_{\mathscr{D}}}\right\rceil<\eta$ (51)
Thus, according to (51) we conclude that $t\eta_{1}+t\eta=t(1+\eta)$. Then, for $t>2$ Sigma index satisfy $\sigma(T)\geqslant t(1+\eta)$. Thus, according to (51) the relationship (50) holds.
Case 2:
If $n$ is odd, where $n=2k+1$. Then, we have according to (49) the relationship should be
$\sigma(T)\geqslant t\eta_{1}\left\lfloor\frac{2k+1}{2k+1-\eta}\right\rfloor+\frac{1}{t}\eta_{1}\left\lceil\frac{2k+1}{\eta-\lambda_{\mathscr{D}}}\right\rceil,$ (52)
Then, we noticed that
$t\eta_{1}\left\lfloor\frac{2k+1}{2k+1-\eta}\right\rfloor+\frac{1}{t}\eta_{1}\left\lceil\frac{2k+1}{\eta-\lambda_{\mathscr{D}}}\right\rceil\leqslant t\eta_{1}(\frac{2k+1}{2k+1-\eta})+\frac{\eta_{1}}{t}(\frac{2k+1}{\eta-\lambda_{\mathscr{D}}})+2\delta.$ (53)
Thus, from (53) we conclude that for $\delta>1,$
$t\eta_{1}(\frac{2k+1}{2k+1-\eta})+\frac{\eta_{1}}{t}(\frac{2k+1}{\eta-\lambda_{\mathscr{D}}})+2\delta\leqslant\eta_{1}(2k+1)\frac{t^{2}(\eta-\lambda_{\mathscr{D}})+(2k-\eta)}{t(2k-\eta)(\eta-\lambda_{\mathscr{D}})}+2(\delta+2).$ (54)
Thus, according to (53) we conclude that $t\eta_{1}+t\eta=t\eta_{1}(2k+1)+2(t\delta+2)$. Then, for $t>2$ and $3\eta>2k$ Sigma index satisfy $\sigma(T)\geqslant t\eta_{1}(2k+1)+2(t\delta+2)$. Thus, according to (53) and (54) the relationship (52) holds.
Therefore, according to to both cases 1 and 2 and according to Proposition 3.4 we find that $\sigma(T)\leqslant\sum_{v\in V(T)}\deg_{T}(v)^{3}$. Thus, the relationship (48) holds for the upper bound of Sigma index. ∎
To clarify the interconnected relationships presented through Theorem 4.5, it is more appropriate to analyze the data in Table 2 and support that analysis with a figure that confirms these relationships.
$\mathscr{D}$ n $\operatorname{irr}(T)$ $\sigma(T)$ $\lambda_{\mathscr{D}}$ $\eta$ $\eta_{1}$
(3,6,8,10,14,16,20) 77 980 16209 11 21 2.12
(7,9,12,14,20,24,27) 113 2050 46312 16.14 31 2.18
(11,12,16,23,26,32,34) 154 3732 107753 22 41 1
(15,15,20,32,32,40,45) 199 6296 233350 28.42 51 3.1
Table 2: Recognize value among Theorem 4.5.
The positions labeled $n,\eta,\eta_{1},\sigma$ and $\operatorname{irr}(T)$ correspond to the variables whose pairwise correlation coefficients are displayed in the correlation matrix as
$M_{\mathscr{D}}=\begin{pmatrix}1.00000&0.998777&0.314881&0.969896&0.990360\\ 0.998777&1.00000&0.282990&0.957017&0.982405\\ 0.314881&0.282990&1.00000&0.497029&0.415811\\ 0.969896&0.957017&0.497029&1.00000&0.994206\\ 0.990360&0.982405&0.415811&0.994206&1.00000\end{pmatrix}$
Noticed that, regression coefficients [1402.91893491,72.8168638][\,-1402.91893491\,,72.8168638\,], intercept $53643.804983904585$, Model $R^{2}$ is $0.9997469414194662$ and predicted Sigma for $n=400$ is 492960.5317043649-492960.5317043649.
Through the following Figure 3, we present analyze the data among Table 2 to consider the results in Theorem 4.5 cleared.
Figure 3: Analysis some of data provided by Table 2.
5 Conclusion
Through this paper, we presented study of bounds on the Sigma and Albertson indices. The results, such as those in Proposition 3.1 with $0<\frac{2\operatorname{irr}_{\min}}{\Delta(\Delta-1)^{2}}<1$ for the Albertson index and in Theorem 4.2 with $\sigma_{\max}(T)\leqslant\dfrac{1}{2(\Delta-3)}\left(\left\lfloor\dfrac{3n^{2}}{4}\right\rfloor\left\lceil\dfrac{n^{2}}{4}\right\rceil\right)$ for the Sigma index, highlight that these indices exhibit significantly larger values for trees compared to connected graphs, as illustrated in Figure 1.
These bounds, derived from degree sequences $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$, maximum degrees $\Delta_{\mathscr{A}}$, $\Delta_{\mathscr{R}}$, and parameters like $\lambda_{\mathscr{D}}$ and $\eta$, provide precise constraints that reflect the interplay between graph topology and index behavior, facilitating applications in extremal graph theory and molecular property prediction.
Furthermore, the established bounds, such as those in Lemma 3.2 with $\sigma(T)\geqslant\sqrt{\lambda_{\mathscr{D}}\sum_{i=1}^{n}d_{i}^{3}}-\left(\left\lfloor\frac{2n}{\lambda_{\mathscr{A}}}\right\rfloor+\left\lceil\frac{2m}{\lambda_{\mathscr{R}}}\right\rceil\right)+(n-\Delta)^{2}$ and Corollary 4.4 with $\sigma(T)>4n-2\eta\lambda_{\mathscr{D}}-(n-\eta)\left\lfloor\frac{n}{n-\eta}\right\rfloor^{2}+(n-\eta)\left\lfloor\frac{n}{n-\lambda_{\mathscr{D}}}\right\rfloor$, emphasize the dependence of the Sigma index on degree heterogeneity and structural parameters. These findings not only advance the theoretical framework for analyzing topological indices but also enhance their utility in quantitative structure-property relationships (QSPR) and network science. By identifying extremal cases and sharp bounds, as seen in Theorem 4.3 and Theorem 4.5, this study provides a robust foundation for optimizing graph-based models and understanding the structural limits of molecular and network systems.
Declarations
  • Funding: Not Funding.
  • Conflict of interest/Competing interests: The author declare that there are no conflicts of interest or competing interests related to this study.
  • Ethics approval and consent to participate: The author contributed equally to this work.
  • Data availability statement: All data is included within the manuscript.